OpenStudy (et365):

Simplify and write the trigonometric expression in terms of sine and cosine: tan^2x-sec^2x?

OpenStudy (campbell_st):

well it can be written as \[\frac{\sin^2(x)}{\cos^2(x)} - \frac{1}{\cos^2(x)} = \frac{\sin^2(x) - 1}{\cos^2(x)}\] not sure how much simplifying is needed, but this is a good start.

OpenStudy (et365):

It wants the equation in a number form from what it looks like.

OpenStudy (campbell_st):

ok \[\sin^2(x) + \cos^2(x) = 1 ~~~ then \cos^2(x) = 1 - \sin^2(x)~~~~ or~~~~-\cos^2(x) = \sin^2(x) - 1\] make the substitution into the equation n the previous post. This should allow you to get a number answer

OpenStudy (campbell_st):

oops \[-\cos^2(x) = \sin^2(x) - 1\]

OpenStudy (et365):

So how should I continue the equation after the left side has been eliminated? The sin^2(x) should be canceled because it is in the numerator correct?

OpenStudy (campbell_st):

if you make the substitution you get -cos^2(x)/cos^2(x) = -1

OpenStudy (et365):

Ohhh I see substitute it into the end